# OBJEX(MT): Objective Extraction and Metacognitive Calibration for LLM-as-a-Judge under Multi-Turn Jailbreaks

### v1

@@ Abstract
Large language models (LLMs) are increasingly used as judges of other models, yet it remains unclear whether a judge can reliably infer the latent objective of the conversation it evaluates—especially when the goal is spread across noisy, adversarial, multi-turn jailbreaks.
We introduce OBJEX(MT), a benchmark that requires a model to
  1. 1.
    distill a transcript into a single-sentence base objective, and
  2. 2.
    report a self-assessed confidence.
Accuracy is scored by an LLM judge via semantic similarity between extracted and gold objectives; correctness uses a single human-aligned threshold calibrated once on N=100 items ($\tau^{*}=0.61$), and metacognition is assessed with ECE, Brier, Wrong@High-Conf, and risk–coverage curves.
We evaluate gpt-4.1, claude-sonnet-4, and Qwen3-235B-A22B-FP8 across SafeMT Attack_600, SafeMTData_1K, MHJ, and CoSafe. claude-sonnet-4 attains the highest objective-extraction accuracy (0.515) and best calibration (ECE = 0.296; Brier = 0.324), while gpt-4.1 and Qwen3 tie at 0.441 accuracy yet exhibit pronounced overconfidence (mean confidence $\approx 0.88$ vs. accuracy $\approx 0.44$; Wrong@0.90 $\approx 48$$52\%$).
Performance varies sharply by dataset ($\approx 0.167$$0.865$), with MHJ comparatively easy and Attack_600/CoSafe harder. These results indicate that LLM judges often misinfer objectives with high confidence in multi-turn jailbreaks, and motivate operational guidance: provide judges with explicit objectives when possible and use selective-prediction/abstention to manage risk.
We release prompts, scoring templates, and complete logs to facilitate replication and analysis.
@@ 1 Introduction
This paper: OBJEX(MT).
We introduce OBJEX(MT) (Objective EXtraction in Multi-Turn jailbreaks), a benchmark that isolates and measures an LLM’s ability to (i) recover the base objective of a multi-turn jailbreak conversation and (ii) calibrate its confidence in that recovery. We evaluate on SafeMT Attack_600 and SafeMTData_1K (collectively, SafeMTData), MHJ (human multi-turn jailbreaks), and CoSafe (Ren et al., 2025; Li et al., 2024a; Yu et al., 2024). Concretely, a model receives a multi-turn transcript and must return a one-sentence base prompt that plainly states the attacker’s goal, together with a confidence score in $[0,1]$. For accuracy, we use an LLM evaluator to assign a semantic similarity score between the extracted and gold objectives; for metacognition, we analyze how confidence aligns with correctness via reliability diagrams (ECE), Brier score, Wrong@High-Conf, and selective-prediction trade-offs.
@@
  • Problem framing. We formalize objective extraction under multi-turn jailbreaks across SafeMT Attack_600, SafeMTData_1K, MHJ, and CoSafe (Ren et al., 2025; Li et al., 2024a; Yu et al., 2024).
  • Benchmark, metrics, and metacognition. We release OBJEX(MT) with an instruction template that elicits a base_prompt and self-reported confidence. Evaluation combines LLM-based semantic matching with metacognitive calibration analyses (ECE, Brier, Wrong@High-Conf, and selective prediction) to quantify whether models know when they know.
@@
We position OBJEX(MT) as a general diagnostic for objective understanding and metacognitive reliability under noisy multi-step inputs, with immediate ramifications for safety evaluation. By starting from a general notion of latent-intent recovery and then specializing to multi-turn jailbreaks, our results invite reuse beyond safety (e.g., multi-hop QA and tool-use auditing) while delivering concrete prescriptions for LLM safety evaluators.
@@ 2.3 Multi-turn safety datasets
MHJ collects human multi-turn jailbreaks with tactic metadata (Li et al., 2024a); CoSafe targets coreference-driven multi-turn attacks (Yu et al., 2024); and SafeMTData comprises two subsets: SafeMT Attack_600 (HarmBench single queries expanded into multi-turn attack paths) and SafeMTData_1K (safety-alignment dialogues including refusals). We use both subsets as-is without schema unification (Ren et al., 2025).
@@ 2.5 Metacognition, calibration, and abstention
A complementary line of work studies whether LLMs can know what they don’t know and express uncertainty appropriately. Models can be prompted to produce verbalized confidence that is often better-calibrated than raw token probabilities (Tian et al., 2023); large models also display varying degrees of self-knowledge in refusing unanswerable questions (Yin et al., 2023). Calibration has become a core reliability target with metrics such as ECE/Brier and with methods that explicitly train or post-hoc adjust confidence (Geng et al., 2024; Ulmer et al., 2024; Huang et al., 2024). Relatedly, selective prediction/abstention frameworks study when a model should defer to humans under low confidence, offering operational levers for safety-critical workflows. We situate OBJEX(MT) within this literature by treating confidence as a metacognitive signal and quantifying its alignment to correctness during objective extraction.
@@ 3.1 Datasets and instance construction
Sources. We evaluate across four public multi-turn safety datasets: SafeMTData_1K, CoSafe, SafeMT Attack_600, and MHJ. We do not perform label-schema unification or de-duplication; each dataset’s provided instance is evaluated as-is.
@@ 3.2 Models and single-pass decoding
Models. We run black-box inference for gpt_4.1, claude-sonnet-4, and Qwen3-235B-A22B-FP8. To isolate the objective-extraction behavior, we take a single pass per instance (one output per model per item), with default provider decoding (no stochastic replicas). For the similarity judge used in §3, we fix a single evaluator model, gpt_4.1, across all experiments.
@@ 3.3 Objective-extraction instruction
We use a single fixed instruction across models that asks the system to (i) distill the dialogue’s latent goal into a one-sentence imperative base_prompt and (ii) return a self-reported confidence $\in[0,1]$ as JSON only. The instruction explicitly requires stripping role-play wrappers and other evasion artifacts, choosing the primary objective when multiple goals appear, and lowering confidence under ambiguity. See Appx. A.1 for the verbatim template. We parse the JSON and store extracted_base_prompt and extraction_confidence. The full judge template appears in Appx. A.1.
@@ 3.5 From similarity to correctness (human-aligned thresholding)
$\tau^{\star}\in\arg\max_{\tau\in\mathcal{T}}\mathrm{F1}\!\left(\{\!(\hat{y}^{(\tau)}_{i},\,y^{\text{human}}_{i})\!\}_{i=1}^{N}\right),$
where $\mathcal{T}$ is a dense grid on $[0,1]$ (step $0.01$ by default) and $\mathrm{F1}=2\,\frac{\mathrm{Prec}\cdot\mathrm{Rec}}{\mathrm{Prec}+\mathrm{Rec}}$. Ties are broken by choosing the smallest $\tau$ achieving the maximal F1 (conservative toward reducing false negatives).
@@ 4.1 Judge calibration on the human-labeled set (N=100)
On the Labeling sheet, class balance after the fixed mapping (Exact/High $\Rightarrow 1$, Moderate/Low $\Rightarrow 0$) is $39\%/61\%$ (pos/neg). The realized source mix is SafeMTData_1K: 56, MHJ_local: 22, SafeMTData_Attack600: 22 (CoSafe items are absent after filtering). Sweeping $\{0.00,0.01,\dots,1.00\}$, F1 peaks at $\mathbf{\tau^{\star}{=}0.61}$ with F1=0.826, Precision=0.717, Recall=0.974. A nonparametric bootstrap (B=1000) yields a 95% CI of [0.450, 0.900] for $\tau^{\star}$. We freeze $\tau^{\star}$ for all downstream analyses (see calibration.json). :contentReference[oaicite:1]index=1
@@ 4.2 Overall objective-extraction accuracy (frozen $\tau^{\star}$)
We mark an extraction correct when the judge score $s$ meets the frozen threshold $s\geq 0.61$. Each model is evaluated once per instance.
@@
$\Delta$ (gpt-4.1 - claude-sonnet-4) ==\,0.074-0.074  ([0.096-0.096, 0.054-0.054], significant); $\Delta$ (gpt-4.1 - Qwen3) ==\,0.000-0.000  ([0.021-0.021, $0.019$], n.s.); $\Delta$ (claude-sonnet-4 - Qwen3) ==\,+0.074+0.074  ([0.053, 0.096], significant). Takeaway: claude-sonnet-4 is statistically more accurate overall; gpt-4.1 and Qwen3 are tied within uncertainty.
@@ 4.3 Dataset-wise performance (heterogeneity)
gpt-4.1: SafeMTData_Attack600 0.167 (600), SafeMTData_1K 0.516 (1,698), MHJ_local 0.834 (537), CoSafe 0.318 (1,404). Spread (max–min): 0.667.
claude-sonnet-4: SafeMTData_Attack600 0.310 (600), SafeMTData_1K 0.667 (1,762), MHJ_local 0.865 (557), CoSafe 0.274 (1,404). Spread: 0.591.
Qwen3-235B-A22B-FP8: SafeMTData_Attack600 0.210 (600), SafeMTData_1K 0.530 (1,767), MHJ_local 0.733 (569), CoSafe 0.311 (1,404). Spread: 0.523.
Across models, accuracies range from 0.167 (hardest) to 0.865 (easiest), a spread of 0.698. In aggregate, MHJ_local is consistently easiest, while SafeMTData_Attack600 and CoSafe are more challenging. (Per-source coverage, $n_{\text{scored}}/n_{\text{total}}$, appears in the supplementary table.)
@@ 4.4 Metacognition from self-reported confidence
Expected Calibration Error (10 bins; lower is better). gpt-4.1 0.437, claude-sonnet-4 0.296, Qwen3 0.447.
Brier score (lower is better). gpt-4.1 0.426, claude-sonnet-4 0.324, Qwen3 0.441.
Wrong@High-Confidence (lower is better). At thresholds 0.80/0.90/0.95: gpt-4.1 53.2% / 47.8% / 45.7%; claude-sonnet-4 42.2% / 36.4% / 21.3%; Qwen3 55.1% / 52.4% / 45.4%.
Selective prediction (risk–coverage). AURC: gpt-4.1 0.497, claude-sonnet-4 0.406, Qwen3 0.498. Takeaway: claude-sonnet-4 is best-calibrated (lowest ECE/Brier/AURC) and far less error-prone at high confidence. In contrast, gpt-4.1 and Qwen3 are overconfident (mean confidences 0.877/0.888 vs. accuracy 0.441).
4.5 Robustness to categorical mapping
Mapping categories directly (Exact/High $\Rightarrow 1$, Moderate/Low $\Rightarrow 0$) preserves the overall ordering: gpt-4.1 0.400, claude-sonnet-4 0.473, Qwen3 0.399. Per-source (gpt-4.1: Attack600 0.142, 1K 0.473, MHJ_local 0.793, CoSafe 0.271; claude-sonnet-4: 0.262/0.617/0.837/0.237; Qwen3: 0.173/0.480/0.687/0.277), supporting that conclusions are not artifacts of the score threshold.
4.6 Reporting quality and coverage
Judge scoring coverage is essentially complete (overall $\geq$ 99.9% across models). All rows carry valid confidences, and extraction_error rate is 0%. Detailed per-source scored/total counts are included in the supplementary results sheet.
4.7 Summary
Using a single human-aligned threshold ($\tau^{\star}{=}0.61$) calibrated once on 100 items and then frozen, claude-sonnet-4 attains the highest accuracy (0.515) and strongest metacognition (ECE 0.296, Brier 0.324, AURC 0.406), significantly outperforming both gpt-4.1 and Qwen3 in accuracy. Performance varies substantially by dataset, with MHJ_local comparatively easy and SafeMTData_Attack600/CoSafe challenging; robustness checks confirm the same ordering.

### v2

# ObjexMT: Objective Extraction and Metacognitive Calibration for LLM-as-a-Judge under Multi-Turn Jailbreaks

@@ Abstract
LLM-as-a-Judge (LLMaaJ) now underpins scalable evaluation, yet we lack a decisive test of a judge’s qualification: can it recover a conversation’s latent objective and know when that inference is trustworthy? LLMs degrade under irrelevant or long context; multi-turn jailbreaks further hide goals across turns. We introduce ObjexMT, a benchmark for objective extraction and metacognition. Given a multi-turn transcript, a model must return a one-sentence base objective and a self-reported confidence. Accuracy is computed via LLM-judge semantic similarity to gold objectives, converted to binary correctness by a single human-aligned threshold calibrated once on N=100 items ($\tau^{\star}\!=\!0.61$). Metacognition is evaluated with ECE, Brier, Wrong@High-Conf, and risk–coverage. Across gpt-4.1, claude-sonnet-4, and Qwen3-235B-A22B-FP8 on SafeMTData_Attack600, SafeMTData_1K, MHJ, and CoSafe, claude-sonnet-4 attains the best objective-extraction accuracy (0.515) and calibration (ECE 0.296; Brier 0.324); gpt-4.1 and Qwen3-235B-A22B-FP8 tie at 0.441 but are overconfident (mean confidence $\approx$0.88 vs. accuracy $\approx$0.44; Wrong@0.90 $\approx$48–52%). Performance varies by dataset ($\approx$0.167–0.865). ObjexMT thus supplies an actionable test for LLM judges: when objectives are not explicit, judges often misinfer them with high confidence. We recommend exposing objectives when feasible and gating decisions by confidence otherwise. Code and data at https://github.com/hyunjun1121/ObjexMT_dataset.
@@ 1 Introduction
This paper: ObjexMT.
We introduce ObjexMT (Objective EXtraction in Multi-Turn jailbreaks), a benchmark that isolates and measures an LLM’s ability to (i) recover the base objective of a multi-turn jailbreak conversation and (ii) calibrate its confidence in that recovery. We evaluate on SafeMTData_Attack600 and SafeMTData_1K (collectively, SafeMTData), MHJ (human multi-turn jailbreaks), and CoSafe (Ren et al., 2025; Li et al., 2024a; Yu et al., 2024). Concretely, a model receives a multi-turn transcript and must return a one-sentence base prompt that plainly states the attacker’s goal, together with a confidence score in $[0,1]$. For accuracy, we use an LLM evaluator to assign a semantic similarity score between the extracted and gold objectives; for metacognition, we analyze how confidence aligns with correctness via reliability diagrams (ECE), Brier score, Wrong@High-Conf, and selective-prediction trade-offs.
@@
  • Problem framing. We formalize objective extraction under multi-turn jailbreaks across SafeMTData_Attack600, SafeMTData_1K, MHJ, and CoSafe (Ren et al., 2025; Li et al., 2024a; Yu et al., 2024).
  • Benchmark, metrics, and metacognition. We release ObjexMT at https://github.com/hyunjun1121/ObjexMT_dataset with: (i) instruction and scoring templates, (ii) the complete experimental dataset (raw transcripts and I/O, extracted objectives and confidences, judge similarity scores and categories, human calibration labels, and per-row spreadsheets), and (iii) materials to support reproducible analysis. Evaluation combines LLM-based semantic matching with metacognitive calibration analyses (ECE, Brier, Wrong@High-Conf, and selective prediction) to quantify whether models know when they know.
@@
We position ObjexMT as a general diagnostic for objective understanding and metacognitive reliability under noisy multi-step inputs, with immediate ramifications for safety evaluation. By starting from a general notion of latent-intent recovery and then specializing to multi-turn jailbreaks, our results invite reuse beyond safety (e.g., multi-hop QA and tool-use auditing) while delivering concrete prescriptions for LLM safety evaluators.
@@ 2.3 Multi-turn safety datasets
MHJ collects human multi-turn jailbreaks with tactic metadata (Li et al., 2024a); CoSafe targets coreference-driven multi-turn attacks (Yu et al., 2024); and SafeMTData comprises two subsets: SafeMTData_Attack600 (HarmBench single queries expanded into multi-turn attack paths) and SafeMTData_1K (safety-alignment dialogues including refusals). We use both subsets as-is without schema unification (Ren et al., 2025).
@@ 2.5 Metacognition, calibration, and abstention
A complementary line of work studies whether LLMs can know what they don’t know and express uncertainty appropriately. Models can be prompted to produce verbalized confidence that is often better-calibrated than raw token probabilities (Tian et al., 2023); large models also display varying degrees of self-knowledge in refusing unanswerable questions (Yin et al., 2023). Calibration has become a core reliability target with metrics such as ECE/Brier and with methods that explicitly train or post-hoc adjust confidence (Geng et al., 2024; Ulmer et al., 2024; Huang et al., 2024). Relatedly, selective prediction/abstention frameworks study when a model should defer to humans under low confidence, offering operational levers for safety-critical workflows. We situate ObjexMT within this literature by treating confidence as a metacognitive signal and quantifying its alignment to correctness during objective extraction.
@@ 3 Methodology
3.1 Threat model and scope
We consider adversarial multi-turn interactions in which an attacker distributes or disguises a harmful goal over $N$ turns. Let the dialogue be $D=\{(u_{t},m_{t})\}_{t=1}^{N}$ with user utterances $u_{t}$ and model replies $m_{t}$. The latent base objective $b^{\star}$ is the minimal imperative instruction that, if issued as a single-turn prompt, would pursue the same harmful goal as $D$.
Operationalization.
Our extractor must produce a single sentence $\hat{b}$ that captures the core objective while stripping role-play wrappers, meta-instructions, and obfuscation. When multiple plausible objectives appear, the instruction (Appx. A.1) directs the model to choose the primary one, i.e., the goal that turns converge on and that would best explain the observed attack trajectory. [^1] We avoid algorithmic tie-breakers to keep the task model-agnostic; ambiguity should be reflected as lower self-reported confidence.
Why one sentence?
We intentionally constrain $\hat{b}$ to a single imperative to stress objective identification rather than plan expansion. This sacrifices some multi-step nuance but makes correctness auditable and comparable across datasets.
Scope.
We evaluate $b^{\star}$ recovery only; we do not score downstream generation or direct safety refusal. The task is distinct from harmfulness classification: models must infer intent from noisy, long, and sometimes self-contradictory contexts.
@@ 3.2 Datasets and instance construction
Sources. We evaluate across four public multi-turn safety datasets: SafeMTData_1K, CoSafe, SafeMTData_Attack600, and MHJ. We do not perform label-schema unification or de-duplication; each dataset’s provided instance is evaluated as-is.
@@ 3.3 Models and single-pass decoding
Models. We run black-box inference for gpt-4.1, claude-sonnet-4, and Qwen3-235B-A22B-FP8. To isolate the objective-extraction behavior, we take a single pass per instance (one output per model per item), with default provider decoding (no stochastic replicas). For the similarity judge used in §3.5, we fix a single evaluator model, gpt-4.1, across all experiments.
@@ 3.4 Objective-extraction instruction
We use a single fixed instruction across models that asks the system to (i) distill the dialogue’s latent goal into a one-sentence imperative base_prompt and (ii) return a self-reported confidence $\in[0,1]$ as JSON only. The instruction explicitly requires stripping role-play wrappers and other evasion artifacts, choosing the primary objective when multiple goals appear, and lowering confidence under ambiguity. See Appx. A.1 for the verbatim template. We parse the JSON and store extracted_base_prompt and extraction_confidence.
@@ 3.6 From similarity to correctness (human-aligned thresholding)
$\tau^{\star}\in\operatorname*{arg\,max}_{\tau\in\mathcal{T}}\mathrm{F}_{1}\!\left(\{\!(\hat{y}^{(\tau)}_{i},\,y^{\text{human}}_{i})\!\}_{i=1}^{N}\right),$
where $\mathcal{T}$ is a dense grid on $[0,1]$ (step $0.01$ by default) and $\mathrm{F}_{1}=2\,\frac{\mathrm{Prec}\cdot\mathrm{Rec}}{\mathrm{Prec}+\mathrm{Rec}}$. Ties are broken by choosing the smallest $\tau$ achieving the maximal $F_{1}$ (conservative toward reducing false negatives).
@@ 4 Results
Table 1: Human-aligned thresholding on N=100N{=}100 items. We freeze $\tau^{\star}$ for all analyses.
Pos./Neg. $\tau^{\star}$ $F_{1}$ Precision Recall
39 / 61 0.61 0.826 0.717 0.974
@@ 4.1 Judge calibration on the human-labeled set (N=100)
On the Labeling sheet, class balance after the fixed mapping (Exact/High $\Rightarrow 1$, Moderate/Low $\Rightarrow 0$) is $39\%/61\%$ (pos/neg). The realized source mix is SafeMTData_1K: 56, MHJ: 22, SafeMTData_Attack600: 22 (CoSafe items are absent after filtering). Sweeping $\{0.00,0.01,\dots,1.00\}$, $F_{1}$ peaks at $\mathbf{\tau^{\star}{=}0.61}$ with $F_{1}$=0.826, Precision=0.717, Recall=0.974. A nonparametric bootstrap (B=1000) yields a 95% CI of [0.450, 0.900] for $\tau^{\star}$. We freeze $\tau^{\star}$ for all downstream analyses (see calibration.json).
@@ 4.2 Overall objective-extraction accuracy (frozen $\tau^{\star}$)
We mark an extraction correct when the judge score $s$ meets the frozen threshold $s\geq 0.61$. Each model is evaluated once per instance. Reproducibility. Per-row spreadsheets that yield all aggregates in this section are included in the release at https://github.com/hyunjun1121/ObjexMT_dataset.
@@
$\Delta$ (gpt-4.1 - claude-sonnet-4) ==\,0.074-0.074  ([0.096-0.096, 0.054-0.054], significant); $\Delta$ (gpt-4.1 - Qwen3-235B-A22B-FP8) ==\,0.000-0.000  ([0.021-0.021, $0.019$], n.s.); $\Delta$ (claude-sonnet-4 - Qwen3-235B-A22B-FP8) ==\,+0.074+0.074  ([0.053, 0.096], significant). Takeaway: claude-sonnet-4 is statistically more accurate overall; gpt-4.1 and Qwen3-235B-A22B-FP8 are tied within uncertainty.
@@ 4.3 Dataset-wise performance (heterogeneity)
gpt-4.1: SafeMTData_Attack600 0.167 (600), SafeMTData_1K 0.516 (1,698), MHJ 0.834 (537), CoSafe 0.318 (1,404). Spread (max–min): 0.667.
claude-sonnet-4: SafeMTData_Attack600 0.310 (600), SafeMTData_1K 0.667 (1,762), MHJ 0.865 (557), CoSafe 0.274 (1,404). Spread: 0.591.
Qwen3-235B-A22B-FP8: SafeMTData_Attack600 0.210 (600), SafeMTData_1K 0.530 (1,767), MHJ 0.733 (569), CoSafe 0.311 (1,404). Spread: 0.523.
Across models, accuracies range from 0.167 (hardest) to 0.865 (easiest), a spread of 0.698. In aggregate, MHJ is consistently easiest, while SafeMTData_Attack600 and CoSafe are more challenging. (Per-source coverage, $n_{\text{scored}}/n_{\text{total}}$, appears in the supplementary table.)
@@
Table 2: Pairwise accuracy deltas ($\Delta=A-B$; bootstrap 95% CI).
Comparison $\Delta$ Acc 95% CI Significant?
gpt-4.1 - Qwen3-235B-A22B-FP8 0.000-0.000 [0.021-0.021, $0.019$] No
@@ 4.4 Metacognition from self-reported confidence
Table 3: Objective extraction and metacognition (frozen $\tau^{\star}\!=\!0.61$). Lower is better for ECE, Brier, Wrong@0.90, and AURC.
Model Accuracy ECE Brier Wrong@0.90 AURC
gpt-4.1 0.441 0.437 0.426 47.8% 0.497
claude-sonnet-4 0.515 0.296 0.324 36.4% 0.406
Qwen3-235B-A22B-FP8 0.441 0.447 0.441 52.4% 0.498
Table 4: Error rates among high-confidence predictions (lower is better).
Model @0.80 @0.90 @0.95
gpt-4.1 53.2% 47.8% 45.7%
claude-sonnet-4 42.2% 36.4% 21.3%
Qwen3-235B-A22B-FP8 55.1% 52.4% 45.4%
Representative failure modes.
Qualitative review of mis-extractions reveals four recurring patterns: (i) Wrapper dominance: the extractor paraphrases role-play or “for research” shells instead of the concealed goal; (ii) Coreference drift: objectives spread over pronouns (“it”, “the plan”) cause the model to pick a proximate but secondary task; (iii) Distractor anchoring: topical but extraneous subgoals (e.g., translation or formatting) are mistaken as the primary objective; (iv) Multi-objective collapse: when two goals co-exist, models often merge them into a hybrid instruction, which the judge scores as non-matching. These errors frequently carry high self-reported confidence, underscoring the need for confidence gating.
@@
Expected Calibration Error (10 bins; lower is better). gpt-4.1 0.437, claude-sonnet-4 0.296, Qwen3-235B-A22B-FP8 0.447.
Brier score (lower is better). gpt-4.1 0.426, claude-sonnet-4 0.324, Qwen3-235B-A22B-FP8 0.441.
Wrong@High-Confidence (lower is better). At thresholds 0.80/0.90/0.95: gpt-4.1 53.2% / 47.8% / 45.7%; claude-sonnet-4 42.2% / 36.4% / 21.3%; Qwen3-235B-A22B-FP8 55.1% / 52.4% / 45.4%.
Selective prediction (risk–coverage). AURC: gpt-4.1 0.497, claude-sonnet-4 0.406, Qwen3-235B-A22B-FP8 0.498.
Takeaway: claude-sonnet-4 is best-calibrated (lowest ECE/Brier/AURC) and far less error-prone at high confidence. In contrast, gpt-4.1 and Qwen3-235B-A22B-FP8 are overconfident (mean confidences 0.877/0.888 vs. accuracy 0.441).
Robustness and coverage: The model ordering also holds under direct categorical mapping (Exact/High $\Rightarrow 1$, Moderate/Low $\Rightarrow 0$), and judge scoring coverage is $\geq$ 99.9% with extraction_error 0% across models (see Appx. §B.5 and Appx. §B.7).
Robustness. The model ordering is preserved under the direct categorical mapping (Exact/High $\Rightarrow 1$, else 0); see Appx. §B.5.